OpenStudy (anonymous):

I'm stuck on how to do this. I know how to use the substitution method, but I've never seen a problem with a number before the y! x+3y=-3 and 9y= -3x-9. Can anyone show me how to at least start this problem? Thanks!

OpenStudy (texaschic101):

x + 3y = -3 x = -3y - 3 you with me so far ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

you subtracted the 3y and added it to the other side

OpenStudy (texaschic101):

now we will sub -3y - 3 in for x in the other equation 9y = -3x - 9 9y = -3(-3y - 3) - 9 --- distribute through the parenthesis 9y = 9y + 9 - 9 -- subtract 9y from both sides 9y - 9y = 9 - 9 0 = 0 if your answer comes out equal, with no variables, then there is INFINITE SOLUTIONS

OpenStudy (anonymous):

Thank you for explaining! The 9 in front of the y was throwing me off. I really appreciate your help :)

OpenStudy (texaschic101):

anytime :)

OpenStudy (texaschic101):

you just mainly want to isolate a variable and then sub that in. it was easy to isolate x because it was by itself.

OpenStudy (texaschic101):

9y = -3x - 9 -- divide by 9 y = -3/9x - 9/9 y = -1/3x - 1 you could have also subbed in -1/3x - 1 in for y to find x, but you would have to deal with fractions then.

OpenStudy (anonymous):

so I do the same thing for every problem? Like if I have x+y=-3 I'd change it to x=-3-y?

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

or y = -3 - x

OpenStudy (anonymous):

Awesome, sounds easy enough :)

OpenStudy (texaschic101):

you will get the hang of it, then you will be helping others to understand it.

OpenStudy (anonymous):

hopefully :) thanks again for your help

OpenStudy (texaschic101):

anytime :)